Solve the following system of linear equations using either the substitution method or the addition method:
1. 2x+y=5
2. 4x+3y=14
Please help and show me the steps so I can learn. Thank you.

Answer by rothauserc(4718)

The problem states
Solve the following system of linear equations using either the substitution method or the addition method:
1. 2x+y=5
2. 4x+3y=14
using equation 1, we have y= 5-2x
substitute for y in equation 2 and we get
4x + 3*(5-2x) = 14
4x + 15-6x = 14
-2x = -1 then
x = 1/2 and 2*1/2 + y = 5 and y = 4